OpenStudy (anonymous):

prove the identity \[2\cos 2\theta+1=\sin \theta \]

OpenStudy (misty1212):

?

OpenStudy (misty1212):

really not clear maybe the equation editor would help

OpenStudy (anonymous):

\[2\cos 2\Theta +1=\sin \Theta\]

OpenStudy (misty1212):

that would be pretty much of a miracle

OpenStudy (xapproachesinfinity):

are you sure that is how it is written \[\]

OpenStudy (anonymous):

no i mean like prove how its true

OpenStudy (misty1212):

since for example if \(\theta=0\) the left side is 3 and the right side is 0

OpenStudy (xapproachesinfinity):

something is wrong with your equation mu friend

OpenStudy (anonymous):

not solve it i.e. 1/cos=sec ;1/cos=1/cos

OpenStudy (xapproachesinfinity):

an identity has to be true for any x( or other variable you are using)

OpenStudy (anonymous):

it is not an identity in fact it may never be true

OpenStudy (lurker):

:>

OpenStudy (xapproachesinfinity):

you cannot prove something that is not an identity

OpenStudy (xapproachesinfinity):

the other question you asked has the same problem

Nnesha (nnesha):

\[2\cos^2 \theta +1 \] ??

OpenStudy (xapproachesinfinity):

when you say prove we expecting that it is true and we want a way to show it is so careful about the question you are asking

OpenStudy (anonymous):

its already true we just need to show it

OpenStudy (xapproachesinfinity):

what you wrote cannot be achieved in any way using those identities

OpenStudy (anonymous):

oh it is true somewhere i lied

OpenStudy (xapproachesinfinity):

where is it true?

OpenStudy (anonymous):

explain please

OpenStudy (xapproachesinfinity):

again if it true for some values, it does not mean it is identity

OpenStudy (anonymous):

good luck solving that equation...

OpenStudy (xapproachesinfinity):

eh that is really beyond the question @satellite73

OpenStudy (xapproachesinfinity):

hahaha

OpenStudy (anonymous):

.-. its on the worksheet itself it says demonstrate that each identity is true

OpenStudy (xapproachesinfinity):

well, the question is wrong itself and good luck in answering it :)

OpenStudy (xapproachesinfinity):

if you did "prove it" we would love to see how

OpenStudy (anonymous):

maybe there is a typo in the question

OpenStudy (xapproachesinfinity):

the most basic approach to check is what @misty1212 tried from the beginning when x=0 you will get 3=0 ? and decide for yourself does that make sense